Write a system of linear equation representing any physical phenomena

(i) Solve the system by analytical method.

(ii) Solve the system by numerical method.



1
Expert's answer
2021-03-18T04:46:42-0400

Suppose two high school kids during lunch break went to get some snacks. The first kid purchased two doughnuts and three sausage rolls and all cost $12. The other kid got two sausage rolls and a doughnut for $7. Assuming that all the snacks of equal size and amount, how much is the cost of a doughnut and a cost of sausage roll?


We can interpret the above statement as follows:


Let "x" represents doughnut

Let "y" represents sausage roll


Then for the first kid,


"2x+3y =12 \qquad \cdots(i)"

For the other kid,

"2y+x= 7 \implies x+2y=7 \qquad \cdots (ii)"

The above is, thus, a system of linear equation with two unknown variable.


(i) Solve the above system by analytical method.


From (ii),


"x=7-2y \qquad \cdots(iii)"

Substitute (iii) in (i)


"2(7-2y)+3y =12\\
14-4y+3y=12\\
14-y =12\\
-y=12-14\\
-y =-2\\
y =2"

Substitute "y=2" in (iii)


"x=7-2y\\
x=7-2(2)\\
x=7-4\\
x=3"

Therefore, the cost of a doughnut is $3 while the cost of a sausage roll is $2.


(ii) Solve the above system by numerical method.


Using Gauss-Jordan Elimination Method,


The augmented matrix for the above system of equation is:


"\left( \begin{array}{cc|c}2&3&12 \\\\ 1&2&7\end{array}\right)"

We'll therefore make zeros in column 1 except the entry at row 1, column 1 (pivot entry).


Divide row 1 by 2, that is "\left(R_1=\frac{R_1}{2}\right)" , we have:


"\left[ \begin{array}{cc|c} 1 & \frac{3}{2} & 6 \\\\ 1 & 2 & 7 \end{array} \right]"



Subtract row 1 from row 2, that is, "\left(R_2=R_2-R_1\right)" , we have:


"\left[ \begin{array}{cc|c} 1 & \frac{3}{2} & 6 \\\\ 0 & \frac{1}{2} & 1 \end{array} \right]"

Make zeros in column 2 except the entry at row 2, column 2 (pivot entry).


Subtract row 2 multiplied by 3 from row 1, that is: "\left(R_1=R_1-\left(3\right)R_2\right)" . We have:


"\left[ \begin{array}{cc|c} 1 & 0 & 3 \\\\ 0 & \frac{1}{2} & 1 \end{array} \right]"



Multiply row 2 by 2 , that is, "\left(R_2=\left(2\right)R_2\right)" , we have:


"\left[ \begin{array}{cc|c} 1 & 0 & 3 \\\\ 0 & 1 & 2 \end{array} \right]"

From the above,


"x=3, y=2"

Therefore, the cost of a doughnut is $3 while the cost of a sausage roll is $2.